A man's age is three times his son's age.Ten years ago he was five times his son's age. By forming two equations and solving them simultaneously,find both of their ages.
Answer by ankor@dixie-net.com(22740)

A man's age is three times his son's age.Ten years ago he was five times his son's age.
By forming two equations and solving them simultaneously,find both of their ages.
:
Let m = man's present age
let s = son's present age
:
Write an equation for each statement:
:
"A man's age is three times his son's age"
m = 3s
m - 3s = 0
:
"Ten years ago he was five times his son's age."
m - 10 = 5(s-10)
m - 10 = 5s - 50
m - 5s = -50 + 10
m - 5s = -40
multiply by -1 to change the signs, add to the first equation
-m + 5s = 40
+m - 3s = 0
------------adding eliminates m, find s
2s = 40
s = 40/2
s = 20 yr is the son's age
then
m = 3(20)
m = 60 is the man's age
:
:
Check this in the statement:
"Ten years ago he was five times his son's age."
60 - 10 = 5(20-10)
50 = 5(10); confirms out solutions